# Embodying Control in Soft Multistable Grippers from morphofunctional co-design

### v1

@@ 3 Discussion
In this study, we introduced a modeling framework and inverse design methodology for soft robotics, discretizing their continuous response using multistable structures. We propose a robotic architecture, The Dome Phalanx Finger, that leverages mechanical multistability to discretize the robots’ configuration space into desired kinematics, which serve as control set points. Exploiting these characteristics, we embodied diverse static and dynamic responses, offering an alternative strategy for achieving various robot functionalities guided solely by the designed geometry and material response. Implementing an energy-based modeling approach was central to our methodology, enabling us to predict and program the DPF’s behavior accurately. Through this approach, we established a robust framework for modeling the finger’s static and dynamic responses, with a mechanics-informed parameter discovery procedure that yields high-fidelity results consistent with experimental data. Significantly, this methodology extends beyond the DPF geometry and its lattice representation, as the combination of mechanics knowledge (in this case from shell theory), data, and automatic relevance determination can be further exploited for a wide variety of geometries and reduced-order models. Utilizing the predictive capabilities of our model, we successfully implement the inverse co-design of the DPFs’ morphology and controlled functionality yielding desired target positions, accessible states, and programmable time responses based on the structure geometry and material viscoelasticity. Our results show how these characteristics can be designed to allow for the modulation of grasping force and unit reset without the need for closed-loop control. Our findings highlight the potential gains in performance when morphology and functionality are concurrently optimized (i.e., co-designed), a result only feasible via the simplified dynamics and modeling afforded by our approach. The implications of our results are far-reaching, suggesting a promising path toward developing robust soft robots. By encoding the control within the morphology and response of the system, we embody a form of mechanical intelligence offering a reduction in control complexity and forgoing the need for data-hungry models and expensive electronics. The resulting reduction in system complexity paves the way for more efficient, cost-accessible, sustainable, and adaptable soft robotics applications.
Materials and Methods
@@
Unit cell and DPA Fabrication– The unit cells are 3D printed using fused deposition modeling (FDM) on a Raise3D Pro2 printer, utilizing Ninjatek Ninjaflex TPU filament [51]. Each finger is post-processed to correct manufacturing errors and seal the air chambers. For the DPG, we printed a PLA base that secures each of the four fingers (see Movie 1 for assembly).
@@
Stiffness Measurement tests– The actuators were loaded onto a specially designed test rig and tested using a universal testing machine (3345 Instron) equipped with a 100N load cell (see Fig. S9b and Movie 4). The test rig constrained all degrees of freedom at the finger’s base, with six different positions equally spaced along 90° to ensure that the initial load was always perpendicular to the tested geometry. We measured the force-displacement response five times per sample at a constant velocity of 10 mm/s. Raw data was evaluated and plotted using Python 3.12.
@@
Finite element analysis– The unit cells are modeled in Abaqus using linear elastic material properties and a combination of S3R and S4R shell elements. A dynamic implicit quasi-static approach is utilized to capture the structure’s instabilities. Geometric nonlinear analysis (Nlgeom) is used, and snap-through is triggered using a pressure load on the dome while the edges are pinned. The unit cell is initially modeled in the stress-free state. Simulations were run for different material thicknesses ($t$) and dome heights ($H$).
The full geometry of the DPA is also modeled in Abaqus using linear elastic material properties and C3D10 3D elements. Snap-through is triggered using a pressure load on every inner wall air chamber. As discussed in the analysis, a fixed boundary condition is applied to the first unit. After dome inversion, different relaxing steps are used to achieve the final stable state.
@@
Robot actuation– Each DPF is actuated using a pressure input controlled by a ball valve. Thea, results displayed in Movies 1 - 5 are produced by applying pressure as slowly as possible to observe the multistable phenomenon.
Author Contributions
All authors designed the research. JCO and HM developed the model and derived the constants. JCO conducted FE simulations. JSR and JCO conducted the experimental measurements and model validation. All authors contributed to the writing and revision of the manuscript and supplementary materials at different stages, and AFA provided guidance throughout the research.
4 Acknowledgements
@@ A.1.1 Unit Cell
The system’s bistable constitutive units are composed of a dome-shaped structure encapsulated by a square chamber that allows for pneumatic actuation and reset by applying positive and negative pressure, respectably (see Fig. S1a). Each dome unit can be geometrically tuned to exhibit a bistable behavior [21], psudo-bistable(metastable) behavior where the is a snap-through instability, but the unit returns to its zero energy state[45, 46], and a monostable behavior (see Fig. S1c). These mechanical responses can be adjusted by modifying the height $H$ and thickness $t$ of the unit as shown in Fig. S1b.
Figure S1: Dome Phalanx Finger geometry and bistable unit. a) Unit cell geometry and parameters (Dome unit + air pressure chamber). b) Bistability map dependent on dome thickness ($t$) and dome height ($H$) for actuator unit cell. c) Monostable (i), Metastable (ii), and bistable (iii) dome unit behavior. d) Dome Phalanx finger geometry has two possible, stable states (Initial and fully actuated states).
@@ A.1.2 Dome Phalanx Finger (DPF)
The DPF is multistable soft system with sequential bistable units (Fig. S1d and Fig. S2b). The finger derives its multistable behavior from bistable domed-shaped shell elements (Fig. S1a), and it can be geometrically tuned to reach and retain different final shapes after dome inversion. The DPF performance and stable kinematic configurations can be tuned by changing the dome height ($H$), Unit Cell size (UC), strain limiting thickness ($t_{L}$), air chamber dimensions ($W_{ch}$ and $U_{L}$), and the spacing between adjacent cells ($\text{U}_{\text{sep}}$). Once fully inverted, contact between adjacent units and the strain-limiting layer induces the system’s global curvature. The bistable domes on each finger segment support programmable deflections as the domes’ final positions dictate the global kinematic configuration (see Fig. S2b). The contact between the dome tip and the adjacent unit’s chamber provides an additional interaction, contributing to the final curvature after all units are activated and the pressure is removed. The strain energy stored in the gripper topology in its activated states is the summation of the individual contributions from the inverted domes (see Fig. S2c). The magnitude of the interaction between neighboring domes is also affected by dome height ($H$) and chamber thickness ($t_{c}$). Together, dome height and chamber thickness determine the degree of interaction between neighboring segments after dome inversion and yield the system’s global curvature (Fig. S2b) and the final tip position of the system (Fig. S2c).
@@
Figure S2: Dome Phalanx Finger geometry and behavior. a) Geometrical parameters that determine the final position and strain energy of the DPF. b) Different stable states of the DPF are encoded by tunning the dome height in the unit cell. c) Strain energy (i) and Tip displacement (ii) as a function of the number of active units. Three different sets of points are encoded into the DPF.
A.2 Energy-based model and constant derivation
A.2.1 Spring Lattice array
@@
Predicting the behavior of multistable DPFs requires consideration of both the unit cells’ and the global geometric parameters. By characterizing the contribution of each DPF subcomponent and their interactions, an energy landscape can be built where each minimum corresponds to a stable state of the system. The resulting strain from the energy minimization process dictates the programmed stable shapes with their geometrical and stiffness characteristics. We represent the DPR as a lattice comprising non-linear[47], linear, and torsional springs (Fig. S3a). The springs’ stiffness and connectivity allow us to map local extensions and rotations to defined energy contributions. We use non-linear springs featuring a ground (unstressed) state and an inverted (stressed) to capture the influence of the bistable domes on the system. To appropriately represent the bistable dome behavior, we position the non-linear springs so that the path of extension coincides with the dome’s tip position between the ground and inverted states. Linear springs capture connections between the bistable units modeled as struts. In the DPF’s case, the strain these segments experience is negligible, behaving as a strain-limiting layer. Consequently, we model these connections as nearly rigid. While the axial strains experienced by this strain-limiting layer are negligible, the layer is sufficiently thick to display bending resistance. We capture the influence of this bending stiffness by introducing torsional springs at the nodes coincident with the strain-limiting layer. The angular displacement experienced by the torsional springs is dictated by the change in the angle $\vartheta$ formed by the segments (see Fig. S3a). Minimizing the system’s total energy, $E_{tot}=\sum_{i=1}^{n}E_{L}+E_{NL}+E_{T}$ where $n$ is the number of units, constrains the space of possible interactions for the lattice elements into a discrete set of stable states. The rich configuration space of our DPF requires establishing a rational method for providing the initial guess and initializing the minimization process. To this end, we implement a geometric base model to generate initial guesses for the optimization algorithm so that the obtained states are in the neighborhood of physically feasible configurations with improved computational time. For this model, we define $k_{l}=\frac{\sqrt{3}}{2}Et_{lim}$ and $k_{t}=\frac{2}{\sqrt{3}}\frac{Et_{lim}}{12(1-\nu^{2})}$ Where $E$ is the young modulus, $\nu$ is the Poisson ratio, and $t_{lim}$ is the thickness of the limiting layer shown in Fig. S2a.
A.2.2 Material Visco-elastic response
@@
Material time response is captured using a variation of the Kelvin-Voigt viscoelastic constitutive model[52]. The model consists of two components in parallel: a spring (representing the elastic behavior) and a dashpot (representing the viscous behavior) see Fig. 3a.
@@
Non-linear Spring (Elastic Element): The spring represents the material’s dome behavior, as explained before. The element follows Eq. 1.
Dashpot (Viscous Element): The dashpot represents the material’s viscous behavior, meaning it resists deformation by dissipating energy as heat. The force exerted by the dashpot is proportional to the velocity of deformation. Mathematically, this relationship is described by $F=\eta\frac{dx}{dt}$, where $\eta$ coefficient is obtained by fitting the dynamic behavior to the experiments for different loading conditions. We obtained $\eta=0.2$ s-1 and utilized it for our dynamic analysis.
A.2.3 Parameter Tuning

### v2

@@ 3 Conclusion
In this study, we introduced a modeling framework and inverse design methodology for soft robotics, discretizing their continuous response using multistable structures. We propose a robotic architecture, The Dome Phalanx Finger, that leverages mechanical multistability to discretize the robots’ configuration space into desired kinematics, which serves as control set points. Exploiting these characteristics, we embodied diverse static and dynamic responses providing an alternative open-loop control strategy for realizing various robotic tasks normally requiring closed-loop approaches guided solely by the designed geometry and material response. Implementing an energy-based modeling approach was central to our methodology, enabling us to accurately predict and program the DPF’s behavior. Through this approach, we established a robust framework for modeling the finger’s static and dynamic responses, with a mechanics-informed parameter discovery procedure that yields high-fidelity results consistent with experimental data. Significantly, this methodology extends beyond the DPF geometry and its lattice representation, as the combination of mechanics knowledge (in this case from shell theory), data, and recursive feature elimination can be further exploited for a wide variety of geometries and reduced-order models. Utilizing the predictive capabilities of our model, we successfully implement the inverse co-design of the DPFs’ morphology and controlled functionality yielding desired target positions, accessible states, and programmable time responses based on the structure geometry and material viscoelasticity. Our results show how these characteristics can be designed to allow for the modulation of grasping force and unit reset without the need for closed-loop control. Our findings highlight the potential gains in performance when morphology and functionality are concurrently optimized (i.e., co-designed), a result only feasible via the simplified dynamics and modeling afforded by our approach. The implications of our results are far-reaching, suggesting a promising path toward developing robust soft robots. By encoding the control within the morphology and response of the system, we embody a form of mechanical intelligence offering a reduction in control complexity and forgoing the need for data-hungry models and expensive electronics. More importantly, this approach is not only beneficial for systems controlled in open loop. As the robot behavior is transformed into a finite, tractable set of positions or set points, the implementation of closed-loop control strategies is simplified by the tracking of a finite number of states instead of the infinite number inherent to continuous systems. Combining our morphological control strategy with conventional sensors can create a streamlined closed-loop control, where sensors can be utilized to verify if the robot is in a specific state, thus reducing the need for constant feedback from the system. The resulting reduction in system complexity paves the way for more efficient, cost-accessible, sustainable, and adaptable soft robotics applications.
4 Experimental Section
@@
Unit cell and DPF Fabrication– The unit cells are 3D printed using fused deposition modeling (FDM) on a Raise3D Pro2 printer, utilizing Ninjatek Ninjaflex TPU filament [52] and Ninjatek Cheetha TPU [53]. Each finger is post-processed to correct manufacturing errors and seal the air chambers. For the DPG, we printed a PLA base that secures each of the four fingers (see Movie 6 for assembly).
@@
Stiffness Measurement tests– The actuators were loaded onto a specially designed test rig and tested using a universal testing machine (3345 Instron) equipped with a 100N load cell (see Figure S16a and Movie 5). The test rig constrained all degrees of freedom at the finger’s base, with six different positions equally spaced along 90° to ensure that the initial load was always perpendicular to the tested geometry. We measured the force-displacement response five times per sample at a constant velocity of 10 mm/s. Raw data was evaluated and plotted using Python 3.12.
@@
Finite element analysis– The unit cells are modeled in Abaqus using linear elastic material properties and a combination of S3R and S4R shell elements. A dynamic implicit quasi-static approach is utilized to capture the structure’s instabilities. Geometric nonlinear analysis (Nlgeom) is used, and snap-through is triggered using a pressure load on the dome while the edges are pinned. The unit cell is initially modeled in the stress-free state. Simulations were run for different material thicknesses ($t$) and dome heights ($H$). The full geometry of the DPF is also modeled in Abaqus using linear elastic material properties and C3D10 3D elements. Snap-through is triggered using a pressure load on every inner wall air chamber. As discussed in the analysis, a fixed boundary condition is applied to the first unit. After dome inversion, different relaxing steps are used to achieve the final stable state (see section A.2 for details).
@@
Robot actuation– Each DPF is actuated using a pressure input controlled by a ball valve. The results displayed in Movies 2 - 3 are produced by applying pressure as slowly as possible to observe the multistable phenomenon.
Pressure and displacement measurement– The dynamic behavior of the DPF is captured in detail using a Photron Fastcam Mini UX100 high-speed camera and the data is processed using Tracker Video Analysis and Modeling Tool. Input pressure is measured using a HONEYWELL ABPDANN010BG2A3 pressure sensor coupled with an Arduino Uno (see section A.7.1 for more details). Each experiment is performed 5 times and the average values are reported.
Acknowledgements
@@ A.1.1 Unit Cell
The system’s bistable constitutive units are composed of a dome-shaped structure encapsulated by a square chamber that allows for pneumatic actuation and reset by applying positive and negative pressure, respectably (see Figure S1a). Each dome unit can be geometrically tuned to exhibit a bistable behavior (Figure S1c (iii)) [23], pseudo-bistable (metastable - Figure S1c (ii)) behavior where the is a snap-through instability, but the unit returns to its zero energy state [47, 48], and a monostable behavior (Figure S1c (i)). These mechanical responses can be adjusted by modifying the height $H$ and thickness $t$ of the unit, as shown in Figure S1b.
Figure S1: Dome unit geometry and mechanics. a) Unit cell geometry and parameters (Dome unit + air pressure chamber). b) Stability map dependent on dome thickness ($t$) and dome height ($H$) for DPF unit cell. c) Monostable (i), Metastable (ii), and Bistable (iii) dome unit behavior. d) Dome Phalanx finger geometry with two possible, stable states (Initial and fully actuated states).
@@ A.1.2 Dome Phalanx Finger (DPF)
The DPF is multistable soft system with sequential dome units (Figure S1d). The finger derives its multistable behavior from domed-shaped shell elements (Figure S1a), and it can be geometrically tuned to reach and retain different final shapes after dome inversion. The DPF performance and stable kinematic configurations can be tuned by changing the dome height ($H$), Unit Cell size (UC), strain limiting thickness ($t_{\text{lim}}$), air channel dimensions ($W_{ch}$ and $t_{\text{mid}}$), spacing between adjacent cells ($\text{U}^{i}_{\text{sep}}$), unit cell length ($\text{U}^{i}_{\text{L}}$), chamber thickness ($t_{ch}$) and dome thickness. Once fully inverted, contact between adjacent units and the strain-limiting layer induces the system’s global curvature. The domes on each finger segment support programmable deflections as the domes’ final positions dictate the global kinematic configuration (see Figure S2b). The contact between the dome tip and the adjacent unit’s chamber provides an additional interaction, contributing to the final curvature after all units are activated and the pressure is removed. The strain energy stored in the gripper topology in its activated states is the summation of the individual contributions from the inverted domes (see Figure S2c). The magnitude of the interaction between neighboring domes is also affected by dome height ($H$) and chamber thickness ($t_{ch}$). Together, dome height and chamber thickness determine the degree of interaction between neighboring segments after dome inversion and yield the system’s global curvature (Figure S2b) and the final tip position of the system (Figure S2c).
@@
Figure S2: Dome Phalanx Finger geometry and behavior. a) Geometrical parameters that determine the final position and strain energy of the DPF. b) Different stable states of the DPF are encoded by tunning the dome height of each unit cell. c) Strain energy (i) and Tip displacement (ii) as a function of the number of active units. Three different sets of points are encoded into the DPF.
Consequently, various dynamic behaviors can be achieved by adjusting the geometrical parameters, thereby enhancing tunability and control. Specifically, the dome height ($H$) and thickness ($t$), along with the limiting layer thickness ($t_{\text{lim}}$), influence the final stable position or set point. Meanwhile, the chamber thickness ($t_{ch}$) and the air channel dimensions ($W_{ch}$ and $t_{\text{mid}}$) govern the overshoot required to reach a target position (see Figure S3). This tunability enables the gripper’s response to be tailored by simply modifying the DPF’s geometric characteristics.
Figure S3: Geometric parameter effect on the time response of DPF. Setpoint, overshoot, and dynamic response can be tuned by combining different parameters. a) Elastic Modulus (E), b) Dome unit Height (H), c) Chamber thickness $t_{ch}$, d) Limiting layer thickness $t_{\text{lim}}$, e) Dome unit thickness ($t$), f) Channel with $W_{\text{ch}}=k_{\text{mid}}UC$
A.2 Finite Element Simulations
Finite element simulations are performed using Abaqus and Python scripting to iterate over different geometrical configurations. Given the dominance of the geometric phenomena in the system, all simulations are done using linear elastic material properties.
A.2.1 Unit Cell Simulation
The unit cells are modeled using S3R and S4R shell elements. The mesh is structured to capture the dome symmetry and its appropriate post-buckling behavior. A dynamic implicit quasi-static approach is utilized to capture the structure’s instabilities. Geometric nonlinear analysis (Nlgeom) is used, and snap-through is triggered using a displacement control method on the dome while the edges are pinned. The unit cell is initially modeled in the stress-free state. Simulations were run for different material thicknesses ($t$), dome heights ($H$), dome radius ($R_{b}$), and Elastic Modulus ($E$). Strain energy vs dome tip displacement is extracted for every step of the simulation (see Figure S4a) which are used to tune the lattice model constant ($k_{b}$ and $\alpha$ in Equation 1).
A.2.2 Dome Unit interaction
The interaction between dome units is modeled by simulating two consecutive units (see Figure S4b). This approach allows us to assess the influence of one unit on the final position of the adjacent dome tip. The interaction is captured using general contact between the surfaces while varying key parameters such as material thickness ($t$), dome height ($H$), dome radius ($r_{b}$), and elastic modulus (E). Tip displacement is recorded over time, and the final displacement values, as a function of the test parameters, are used to calibrate the parameter $d$ (Equation 1) in the lattice model.
A.2.3 DPF 3D Simulation
The complete geometry of the DPF is modeled using C3D10 3D elements to capture the detailed deformation behavior. Snap-through is initiated by applying a pressure load to each inner wall of the air chambers. 3D simulations are chosen to capture the dynamic behavior under internal pressure loading. A fixed boundary condition is applied to the first unit to prevent rigid body motion. After dome inversion, a series of relaxation steps are performed to allow the structure to reach its final stable state (see Figure S4 c). To validate the lattice model, simulations are conducted for various geometric parameters, including the number of units ($N_{\text{Units}}$), material thickness ($t$), chamber wall thickness ($t_{\text{ch}}$), elastic modulus (E), dome height ($H$), and base radius ($r_{b}$). The final deformed shape is extracted for comparison with experimental results. The entire DPF geometry and simulation workflow is implemented using Python scripting in Abaqus, ensuring efficient parameter exploration and model validation.
Figure S4: Finite Element (FE) simulations for model parameter tuning and validation. (a) Dome unit inversion sequence used to determine $k_{b}$ and $\alpha$. (b) Interaction between dome units to determine $d$. (c) 3D simulation for validating the model’s predictions against FE data.
A.3 Energy-based model and constant derivation
A.3.1 Spring Lattice array
@@
Predicting the behavior of multistable DPFs requires consideration of both the unit cells’ and the global geometric parameters. By characterizing the contribution of each DPF subcomponent and their interactions, an energy landscape can be built where each minimum corresponds to a stable state of the system. The resulting strain from the energy minimization process dictates the programmed stable shapes with their geometrical and stiffness characteristics. We represent the DPR as a lattice comprising nonlinear [49], linear, and torsional springs (Figure S5a). The springs’ stiffness and connectivity allow us to map local extensions and rotations to defined energy contributions. We use nonlinear springs featuring a ground (unstressed) state and an inverted (stressed) to capture the influence of the bistable domes on the system. To appropriately represent the bistable dome behavior, we position the nonlinear springs so that the path of extension coincides with the dome’s tip position between the ground and inverted states. Linear springs capture connections between the nonlinear units modeled as struts. In the DPF’s case, the strain these segments experience is negligible, behaving as a strain-limiting layer. Consequently, we model these connections as nearly rigid. While the axial strains experienced by this strain-limiting layer are negligible, the layer is sufficiently thick to display bending resistance. We capture the influence of this bending stiffness by introducing torsional springs at the nodes coincident with the strain-limiting layer. The angular displacement experienced by the torsional springs is dictated by the change in the angle $\vartheta$ formed by the segments (see Figure S5a).We define $k_{l}=\frac{\sqrt{3}}{2}E\frac{t_{\text{lim}}UC}{U_{L}+t_{\text{ch}}+U_{\text{sep}}}$ and $k_{t}=\frac{2}{\sqrt{3}}\frac{E(t_{lim}+t)}{12(1-\nu^{2})}$ Where $E$ is the young modulus, $\nu$ is the Poisson ratio, and $t_{lim}$ is the thickness of the limiting layer shown in Figure S2a.
Minimizing the system’s total energy, $E_{tot}=\sum_{i=1}^{n}E_{L}+E_{NL}+E_{T}$ where $n$ is the number of units, constrains the space of possible interactions for the lattice elements into a discrete set of stable states. The rich configuration space of our DPF requires establishing a rational method for providing the initial guess and initializing the minimization process. To this end, we implement a geometric base model to generate initial guesses for the optimization algorithm so that the obtained states are in the neighborhood of physically feasible configurations with improved computational time. The geometric model for the dome phalanx finger assumes a fixed extension of the nonlinear spring and pure rotation from segment to segment. Here, the activated states of the finger (i.e., with units with inverted domes) are determined by extending the length of the linkage corresponding to the bistable dome structure such that it reflects the distance along the central axis between the tip of the dome to the base of the dome chamber. The now extended segment yields an angle $\vartheta$ between the present and initial position of the front segment (see Figure S5 a). The model uses this $\vartheta$ to calculate the rotation of the subsequent segment, and the process is repeated until the number of segments is exhausted. In this model, the displacement of the overall fingertip can be calculated as a summation of sines and cosines. This purely geometric model does not account for the internal force balancing characteristic of each stable state, which ultimately leads to errors in the configuration predictions. Nevertheless, predictions of a purely geometric model are a powerful tool in finding solutions within the hyperdimensional energy landscape of the spring lattice when we use the outputs of the geometric model to form the initial guess in our search for local extrema. The stable configurations that result from balancing the force and energy contribution of coupled springs differ from the activated configurations determined by the geometric model, but the discrepancy is one easily overcome by the energy minimization process.
A.3.2 Material viscoelastic response
@@
Material time response is captured using a variation of the Kelvin-Voigt viscoelastic constitutive model[54]. The model consists of two components in parallel: a spring (representing the elastic behavior) and a dashpot (representing the viscous behavior) see Figure 3a.
@@
Nonlinear Spring (Elastic Element): The spring represents the material’s dome behavior, as explained in the previous section. This spring follows Equation 1.
Dashpot (Viscous Element): The dashpot represents the material’s viscous behavior, meaning it resists deformation by dissipating energy as heat. This behavior is modeled by considering two different damping forces, isotropic damping, and internal damping [55]. The isotropic damping for numerical stability on a node $i$ can be written as:
$F_{d}^{iso}(\dot{x}_{i})=\eta_{iso}\dot{x}_{i}.$ (S.10)
where $x_{i}$ represents the position and node $i$ and $\dot{x}_{i}$ represents its velocity. Internal damping is selected to be proportional to the spring coefficients as [55]:
$F_{d}^{\text{int }}\left(x_{ij},\dot{x}_{i},\dot{x}_{j}\right)=\eta_{\text{int}}\left(\left(1-\frac{s_{ij}}{\left\|x_{ij}\right\|}\right)\left(\dot{x}_{i}-\dot{x}_{j}\right)+\frac{s_{ij}}{\left\|x_{ij}\right\|^{3}}\left[x_{ij}\cdot\left(\dot{x}_{i}-\dot{x}_{j}\right)\right]x_{ij}\right)$ (S.11)
where $s_{ij}$ is the rest length of the spring, and $x_{ij}=x_{i}-x_{j}$. $\eta_{iso}$ is selected to guarantee numerical stability, and $\eta_{\text{int }}$ coefficient is obtained by fitting the dynamic behavior to the experiments for different loading conditions. We obtained $\eta=0.05$ Ton s-1 by utilizing the Prony series characterized in [56].
A.3.3 Parameter Tuning